# OCRBench v2: An Improved Benchmark for Evaluating Large Multimodal Models on Visual Text Localization and Reasoning

### v1

@@ Appendix G Samples for LMMs’ Limitations
Figure 21 to Figure 23 provide examples corresponding to the findings discussed in Section 4.4 of the main text, which show error results of GPT-4o [1], Monkey [48], and Qwen2VL-8B on various tasks in OCRBench v2. These examples highlight the current limitations of LLMs on OCR tasks. For instance, LLMs exhibit poor recognition of less frequently encountered texts, struggle to accurately locate text in tasks involving text and coordinates, and demonstrate insufficient perception of text in complex layouts such as rotated texts. Additionally, their logical reasoning abilities are limited when addressing mathematical problems, and their analysis of complex elements in charts remains weak. These are the capabilities of LLMs in OCR tasks that require further improvement.

### v2

@@ A.11 Samples for LMMs’ Limitations
Fig. 19 to Fig. 21 provide examples corresponding to the findings discussed in Sec. 5.3 of the main text, which show error results of GPT-4o [1], Monkey [51], and Qwen2VL-8B on various tasks in OCRBench v2. These examples highlight the current limitations of LLMs on OCR tasks. For instance, LLMs exhibit poor recognition of less frequently encountered texts, struggle to accurately locate text in tasks involving text and coordinates, and demonstrate insufficient perception of text in complex layouts such as rotated texts. Additionally, their logical reasoning abilities are limited when addressing mathematical problems, and their analysis of complex elements in charts remains weak. These are the capabilities of LLMs in OCR tasks that require further improvement.
A.12 Broader Impacts
Our benchmark aims to enhance the evaluation of LMMs in text-oriented visual comprehension tasks. By establishing comprehensive benchmarks that reveal deficiencies in models’ OCR capabilities, we provide insights for improving model performance. This advancement will elevate processing efficiency across scenarios such as document automation, assisted reading tools, and complex layout analysis, thereby benefiting applications in domains like healthcare and education. However, enhanced OCR functionality also introduces risks of misuse, including unauthorized extraction of sensitive information from images, surveillance-related applications, or generation of forged documents. To mitigate these risks, we restrict the use of this benchmark solely to research purposes and urge the community to prioritize privacy and fairness considerations in future model development.
A.13 Limitations
One challenge we encountered is that LMMs sometimes produce responses that deviate from the given instructions, making it difficult to extract the desired answers. In future work, we plan to develop a more objective assessment framework to address this issue.
Another limitation arises when evaluating commercial LMMs, as some models occasionally refuse to answer certain questions due to safety filters or unclear content policies. This can lead to incomplete or biased performance assessments compared to open-source models that do not exhibit such behavior.
Figure 10: Samples for each task.
Figure 11: Samples for each task.
Figure 12: Samples for each task.
Figure 13: Samples for each task.
Figure 14: Samples for each task.
Figure 15: Samples for each task.
Figure 16: Samples for each task.
Figure 17: Samples for each task.
Figure 18: Samples for each task.